# Egorov-Type Semiclassical Limits for Open Quantum Systems with a Bi-Lindblad Structure

### v1

@@ 2.4 Symplectization and the contact–homogeneous correspondence
is a symplectization in the sense of [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]: there exists a nowhere-vanishing function $\sigma$ (here $\sigma=r$) such that $\theta=\sigma\,\Sigma^{\ast}\alpha$.
@@
following the conventions in [C.d.L+2023, Lopez-Gordon2024]. This defines a homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$.
The following result summarizes the correspondence between integrability in the contact and homogeneous symplectic settings (see [C.d.L+2023, Lopez-Gordon2024, G.I.M+2004, Petalidou2002] for proofs and variants, and [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] for related Jacobi and Jacobi–Nijenhuis structures).
@@
    The map $f\mapsto f^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\sigma\,\Sigma^{\ast}f$ defines a bijection between $\mathscr{C}^{\infty}(C)$ and the space of $1$-homogeneous functions on $C^{\Sigma}$.
@@
    $\{f^{\Sigma},g^{\Sigma}\}_{\theta}=\{f,g\}_{\alpha}^{\Sigma}.$
@@
The above result is classical. Items (1)–(3) follow from the standard identification between Jacobi manifolds and homogeneous Poisson manifolds under symplectization; see [G.I.M+2004, Petalidou2002, D.L.M1991]. The equivalence of contact integrability and homogeneous integrability in item (4) is established in [C.d.L+2023, Lopez-Gordon2024]. We record the proposition here for completeness and to fix notation for the subsequent sections, where this correspondence becomes the bridge to the quantum (Lindbladian) setting.
Remark 2.10.
@@
In contrast, starting directly from compatible Jacobi structures on $(C,\alpha)$ [NunesdaCosta1998a, M.M.P1999, P.N2003, C.D.N2010] and using the associated Jacobi–Nijenhuis recursion operators does not, in general, lead to a maximal family of dissipated quantities in involution for a contact system, as we shall discuss later in connection with a no-go result of “Jacobi/bi-Hamiltonian” type. This dichotomy will be crucial when we pass from classical homogeneous and contact dynamics to quantum GKSL generators in the subsequent sections.
@@
[…]
@@
We adopt the convention for the contact Hamiltonian vector field $X_{h}$:
@@
For $\alpha=\mathrm{d}z-p\,\mathrm{d}q$ this yields the coordinate expressions
$\dot{q}=-\frac{\partial h}{\partial p},\qquad\dot{p}=\frac{\partial h}{\partial q}+p\frac{\partial h}{\partial z},\qquad\dot{z}=h-p\frac{\partial h}{\partial p}.$
@@
Then
$\frac{\partial h}{\partial p}=-1,\qquad\frac{\partial h}{\partial q}=0,\qquad\frac{\partial h}{\partial z}=1,$
and the resulting contact dynamics is
$\dot{q}=1,\qquad\dot{p}=p,\qquad\dot{z}=z.$
Hence the flow is explicitly solvable:
$q(t)=q_{0}+t,\qquad p(t)=p_{0}e^{t},\qquad z(t)=z_{0}e^{t}.$
Define the observable
$I_{1}(q,p,z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
Since $R(h)=1$, the condition for $I_{1}$ to be a dissipated quantity is
$X_{h}(I_{1})+(Rh)\,I_{1}=0.$
A direct computation gives
$X_{h}(I_{1})=-\,\frac{\partial I_{1}}{\partial z}=-(-e^{-z})=e^{-z},$
and therefore
$X_{h}(I_{1})+(Rh)\,I_{1}=e^{-z}-e^{-z}=0.$
Equivalently, in terms of the Jacobi bracket,
$\{I_{1},h\}_{\alpha}=\frac{\partial I_{1}}{\partial z}-\frac{\partial I_{1}}{\partial p}+I_{1}=(-e^{-z})-0+e^{-z}=0.$
Hence $I_{1}$ is a dissipated quantity in the sense of Definition 2.7.
Note that $I_{1}$ is not constant along the contact flow:
$X_{h}(I_{1})=e^{-z}\neq 0,$
which illustrates that dissipated quantities (Jacobi integrals) need not be conserved along the flow, unless additional conditions such as $R(I_{1})=0$ are satisfied.
Moreover, $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h$ is trivially dissipated since $\{h,h\}_{\alpha}=0$, and we have
$\{I_{0},I_{1}\}_{\alpha}=0.$
The differentials
$\mathrm{d}h=\mathrm{d}z-\mathrm{d}p,\qquad\mathrm{d}I_{1}=-e^{-z}\,\mathrm{d}z$
are linearly independent on a dense open subset of $C$, so
@@
Therefore $(C,\alpha,h)$ is a completely integrable contact Hamiltonian system with $n=1$, admitting the two commuting dissipated quantities $I_{0}=h$ and $I_{1}=e^{-z}$.
Finally, this contact system arises from symplectization with our sign conventions. Let
[…]
and define the $1$-homogeneous Hamiltonian
$H^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\,r\,h=-r(z-p).$
Then $(C^{\Sigma},\theta,H^{\Sigma})$ is an exact symplectic homogeneous Hamiltonian system, and the contact dynamics of $(C,\alpha,h)$ is recovered by restriction to the hypersurface $\{r=1\}$, in accordance with Proposition 2.8.
@@ 3.1 Bi-Hamiltonian structures on Poisson–Lie groups
is Poisson. In this case $(G,\pi_{0})$ is a Poisson–Lie group; see, e.g., [Lichnerowicz1977, Lichnerowicz1977a, Lichnerowicz1978, D.L.M1991, G.U2003].
@@
On a symplectic manifold $(M,\omega)$ we define the Hamiltonian vector field $X_{H}$ by $\iota_{X_{H}}\omega=\mathrm{d}H$.
Definition 3.1.
A bi-Hamiltonian Poisson–Lie system on a Lie group $G$ consists of:
  • two Poisson–Lie structures $\pi_{0},\pi_{1}\in\Gamma(\wedge^{2}{\mathsf{T}}G)$ on $G$;
  • two Hamiltonians $H_{0},H_{1}\in\mathscr{C}^{\infty}(G)$;
such that:
  1. 1.
    $\pi_{0}$ and $\pi_{1}$ are compatible Poisson tensors in the sense that every linear combination
    is a Poisson tensor on $G$;
  2. 2.
    the corresponding Hamiltonian vector fields coincide,
    $X\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\pi_{0}^{\sharp}(\mathrm{d}H_{0})=\pi_{1}^{\sharp}(\mathrm{d}H_{1}).$
Thus $(\pi_{0},\pi_{1})$ defines a Poisson pencil on $G$, and on each symplectic leaf of $\pi_{0}$ (or $\pi_{1}$) the restriction of $(\pi_{0},\pi_{1})$ and $X$ yields an ordinary bi-Hamiltonian system in the sense of Section 2. The general bi-Hamiltonian theory on Poisson manifolds (and in particular on symplectic leaves) is developed in [C.M.P1993, M.C.F+1997, F.M.P2000, Fernandes1994, G.M.S2003].
Many classical integrable systems, such as the Euler top and its deformations, admit realizations of this type on Poisson–Lie groups [M.P1996, NunesdaCosta1998, Petalidou2000]. In what follows, we shall assume that the bi-Hamiltonian dynamics is (Liouville) integrable on a suitable open dense subset of $G$, and concentrate on how this structure lifts to an exact symplectic manifold and then induces a contact integrable system via homogeneity.
@@ 3.2 Exact symplectic realizations and homogeneous lifts
Let $(G,\pi_{0})$ be a Poisson manifold. A symplectic realization of $(G,\pi_{0})$ is a symplectic manifold $(M,\omega)$ together with a Poisson map $J\colon(M,\omega)\to(G,\pi_{0})$ whose image is an open subset of $G$; see, for instance, [L.M1987, D.L.M1991, G.I.M+2004, Lopez-Gordon2024]. We are interested in the case in which $(M,\omega)$ is exact.
@@
Requiring a symplectic realization $(M,\omega)$ to be exact, $\omega=-\mathrm{d}\theta$, imposes topological restrictions on the realizing manifold $M$ (in particular, $M$ cannot be compact and $\omega]=0$ in de Rham cohomology). This requirement, however, does not constrain the underlying Lie–Poisson or Poisson–Lie structure itself. For a large class of Lie–Poisson systems, including linear and solvable cases, canonical exact realizations are provided by cotangent bundles ${\mathsf{T}}^{\ast}Q$, which are always non-compact and carry a natural exact symplectic structure. The example in Section 3.4 illustrates this situation explicitly. More generally, since our construction only requires working on open dense subsets where the homogeneous integrals are independent, the assumption of exactness should be regarded as a natural geometric convenience rather than a substantive restriction of the Poisson–Lie framework.
Let $\Delta$ be the Liouville vector field of $(M,\theta)$, defined by
$\iota_{\Delta}\omega=-\theta.$
As recalled in Section 2, a function $H\in\mathscr{C}^{\infty}(M)$ is said to be $1$-homogeneous if
$\mathcal{L}_{\Delta}H=H,$
and a triple $(M,\theta,H)$ with $H$ 11-homogeneous is a homogeneous Hamiltonian system. We shall also use the notion of homogeneous integrability from Definition 2.3.
@@
  • a $1$-homogeneous Hamiltonian $H\in\mathscr{C}^{\infty}(M)$ with respect to the Liouville vector field $\Delta$ of $(M,\theta)$,
@@
In many examples of interest, the Poisson pencil $(\pi_{0},\pi_{1})$ admits a symplectic realization $(M,\omega)$ on which the pullbacks of $\pi_{0},\pi_{1}$ give rise to a Poisson–Nijenhuis structure $(\omega,N)$ on $M$ [K.M1990, Wade2001, D.M.P2011, B.K.M2022, B.K.M2022a, B.K.M2024b, Konyaev2021]. The eigenvalues of $N$ then provide a natural candidate for homogeneous integrals in involution, and the homogeneity of $H$ can often be arranged by a suitable choice of exact symplectic structure (or by considering an appropriate homogeneous extension).
@@ 3.3 Contact realizations from homogeneous lifts
We first recall a standard fact about exact symplectic manifolds (see, e.g., [L.M1987, D.L.M1991, Lopez-Gordon2024]).
@@
Hence $\Theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta\wedge(\mathrm{d}\theta)^{n}$ is nonvanishing at every point where $\Delta\neq 0$.
@@
Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ and $(M,\theta,H)$ be as in Assumption 3.5. Let $\Delta$ be the Liouville vector field of $(M,\theta)$ and let $C\subset M$ be a hypersurface transverse to $\Delta$, contained in the open set where $H\neq 0$. Define
@@
Since $H$ is $1$-homogeneous with respect to $\Delta$, we have $\mathcal{L}_{\Delta}H=H$. Writing $\varphi_{s}$ for the flow of $\Delta$, this implies
@@
so, on any integral curve of $\Delta$,
@@
In particular, on the open region where $H\neq 0$, the level sets of $H$ are transversal to the Liouville flow, and every Liouville orbit intersects each nonzero level set exactly once. By construction $C$ is a hypersurface transverse to $\Delta$ and contained in $\{H\neq 0\}$, so every Liouville orbit in the region of interest intersects $C$ exactly once.
@@
Standard results on the contactization of homogeneous Hamiltonian systems (see, for instance, [D.L.M1991, I.L.M+1997, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]) ensure that:
@@
Moreover, since the $F_{j}$ are in Poisson-involution on $M$ and the Jacobi bracket on $C$ is obtained by reduction along the Liouville flow (again, see [D.L.M1991, G.I.M+2004, Petalidou2002, Lopez-Gordon2024]), the restrictions $I_{j}$ are in Jacobi-involution:
@@
(ii) Functional independence. Let $U\subset M$ be the dense open subset where $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are independent. Since $C$ is transverse to the Liouville flow, the intersection $U\cap C$ is a dense open subset of $C$. At any $x\in U\cap C$ the differentials $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are linearly independent in ${\mathsf{T}}^{\ast}_{x}M$. Restricting to ${\mathsf{T}}_{x}C\subset{\mathsf{T}}_{x}M$ we obtain the differentials $\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}$, which remain linearly independent on ${\mathsf{T}}_{x}C$: otherwise there would be a nontrivial linear relation among the $\mathrm{d}F_{j}$ on ${\mathsf{T}}_{x}M$. Thus,
@@
The construction above crucially uses Poisson compatibility and homogeneity on the exact symplectic realization $(M,\theta)$, rather than compatibility of Jacobi structures directly on the contact manifold $(C,\alpha)$. As shown in our previous work on the Jacobi/bi-Hamiltonian no-go phenomenon [C.d.L+2023, Lopez-Gordon2024] (building on [NunesdaCosta1998a, M.M.P1999, N.P2002, C.D.N2010, P.N2003, Petalidou2002]), compatible Jacobi pairs on a contact manifold do not provide enough independent eigenvalues of the associated recursion operator to recover a full set of contact integrals. By passing through the homogeneous Poisson picture on $(M,\theta)$ we circumvent this obstruction: the bi-Hamiltonian structure lives in the Poisson category, while the contact dynamics is obtained by projection along the Liouville flow.
@@
Figure 1: From a bi-Hamiltonian Poisson–Lie system on $(G,\pi_{0},\pi_{1})$ to a contact integrable system $(C,\alpha,h)$ via an exact, homogeneous symplectic realization $(M,\theta,H)$ and a Liouville-transverse hypersurface $C$.
@@
This is an immediate consequence of Proposition 3.7, together with the fact that Liouville-transverse hypersurfaces can be chosen inside any given nonzero level set of $H$ in the region where the homogeneous integrals $(F_{0},\dots,F_{n})$ are independent. ∎
3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two nontrivial contact integrals
We illustrate Proposition 3.7 with the Poisson–Nijenhuis example already discussed in Section 2, now interpreted as a homogeneous bi-Hamiltonian system producing a contact integrable system with the correct number of dissipated quantities for the $3$-dimensional ($n=1$) contact setting.

### v2

@@ 2.4 Symplectization and the contact–homogeneous correspondence
is a symplectization in the sense of [13, 35, 30, 55, 43]: there exists a nowhere-vanishing function $\sigma$ (here $\sigma=r$) such that $\theta=\sigma\,\Sigma^{\ast}\alpha$.
@@
following the conventions in [10, 43]. This defines a homogeneous Hamiltonian system $(C^{\Sigma},\theta,H^{\Sigma})$. The Liouville vector field of $\theta$ in this case is given by $\Delta=r\partial_{r}$.
The following result summarizes the correspondence between integrability in the contact and homogeneous symplectic settings (see [10, 43, 30, 55] for proofs and variants)
@@
    The map $f\mapsto f^{\Sigma}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-\sigma\,\Sigma^{\ast}f$ defines a bijection between $\mathscr{C}^{\infty}(C)$ and the space of $1$-homogeneous functions on $C^{\Sigma}$. This bijection has an associated correspondence between the associated Hamiltonian vector fields given by
    $X_{f^{\Sigma}}=-X_{f}+rR(f)\partial_{r}\,$,
    where $X_{f^{\Sigma}}$ (resp. $X_{f}$) denotes the Hamiltonian vector field of $f^{\Sigma}$ (resp. $f$) with respect to $\omega$ (resp. $\alpha$).
@@
    $\{f^{\Sigma},g^{\Sigma}\}_{\theta}=-\{f,g\}_{\alpha}^{\Sigma}.$
@@
In contrast, starting directly from compatible Jacobi structures on $(C,\alpha)$ [53, 47, 52, 9] and using the associated Jacobi–Nijenhuis recursion operators does not lead to a maximal family of dissipated quantities in involution for a contact system. In fact, in [11] we have proven a no-go theorem. This dichotomy will be crucial when we pass from classical homogeneous and contact dynamics to quantum GKSL generators in the subsequent sections.
@@
$\alpha\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathrm{d}z-p\,\mathrm{d}q.$
@@
The contact Hamiltonian vector field $X_{h}$ is determined by
@@
This yields the coordinate expression
$X_{h}=-\frac{\partial h}{\partial p}\frac{\partial}{\partial q}+\left(\frac{\partial h}{\partial q}+p\frac{\partial h}{\partial z}\right)\frac{\partial}{\partial p}+\left(h-p\frac{\partial h}{\partial p}\right)\frac{\partial}{\partial z}\,.$
@@
Then the resulting contact dynamics is
$X_{h}=\frac{\partial}{\partial q}+p\frac{\partial}{\partial p}+z\frac{\partial}{\partial z}\,$,
whose flow $\varphi_{t}$ is given by
$\varphi_{t}\left(q,p,z\right)=\left(q+t,pe^{t},ze^{t}\right)\,.$
Since $R(h)=1$, the condition for a function $I_{i}$ to be a dissipated quantity is
$0=\{h,I_{i}\}=X_{h}I_{i}-(Rh)I_{i}=X_{h}I_{i}-I_{i}=\frac{\partial I_{i}}{\partial q}+p\frac{\partial I_{i}}{\partial p}+z\frac{\partial I_{i}}{\partial z}-I_{i}\,.$
For instance, $I_{1}=p,I_{2}=z$ and $I_{3}=e^{q}$ are dissipated quantities. Note that they are not constant along the flow of $X_{h}$. In fact, $I_{i}\circ\varphi_{t}=e^{t}I_{i},\,i=1,2,3$. This illustrates that dissipated quantities (Jacobi integrals) need not be conserved along the flow, unless additional conditions such as $R(I_{i})=0$ are satisfied.
Moreover, $I_{0}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h$ is trivially dissipated since the Jacobi bracket is skew-symmetric. The differentials
$\mathrm{d}h=\mathrm{d}z-\mathrm{d}p,\qquad\mathrm{d}I_{1}=\mathrm{d}z$
are linearly independent on $C$, so
@@
Therefore $(C,\alpha,h)$ is a completely integrable contact Hamiltonian system with $n=1$, admitting the two commuting dissipated quantities $I_{0}=h$ and $I_{1}=z$.
The associated homogeneous integrable system $(M,\theta,H,F)$ is given by
$M=C\times\mathbb{R}_{>0}\,,\quad\theta=r\alpha=r\mathrm{d}z-rp\mathrm{d}q\,,\quad H=-rh=rp-rz\,,\quad F=-rI_{1}=-rp\,$,
where $r$ is the canonical coordinate of $\mathbb{R}_{>0}$. The Liouville vector field is given by $\Delta=r\partial_{r}$ The Hamiltonian vector field of $H$ with respect to $\omega=-\mathrm{d}\theta$ reads
$X_{H}=-X_{h}+rR(h)\frac{\partial}{\partial r}=\frac{\partial}{\partial q}+p\frac{\partial}{\partial p}+z\frac{\partial}{\partial z}+r\frac{\partial}{\partial r}\,.$
One can verify that it is 0-homogeneous, i.e., $\Delta,X_{H}]=0$.
Remark 2.11.
In order to simplify the exposition, we will consider only co-oriented contact manifolds and their trivial symplectizations. More generally, one can define a contact distribution on a manifold $M$ as a maximally non-integrable sub-bundle $D\subset{\mathsf{T}}M$ of corank $1$. A contact form is a one-form $\alpha$ such that $\ker\alpha=D$, which may only exist locally, and is never unique (as $f\alpha$ is also a contact form for any non-vanishing function $f$). A more general notion of symplectization is an $\mathbb{R}^{\times}$-principal bundle $P\to M$ endowed with a symplectic form which is homogeneous with respect to the principal action. For more details, see [7, 31, 27, 29]
@@ 3.1 Bi-Hamiltonian structures on Poisson–Lie groups
is Poisson. In this case $(G,\pi_{0})$ is a Poisson–Lie group; see [3] and references therein.
@@
Many classical integrable systems, such as the Euler top and its deformations, admit realizations of this type on Poisson–Lie groups [49, 51, 54]. In what follows, we shall assume that the bi-Hamiltonian dynamics is (Liouville) integrable on a suitable open dense subset of $G$, and concentrate on how this structure lifts to an exact symplectic manifold and then induces a contact integrable system via homogeneity.
@@ 3.2 Exact symplectic realizations and homogeneous lifts
Let $(G,\pi_{0})$ be a Poisson manifold. A symplectic realization of $(G,\pi_{0})$ is a symplectic manifold $(M,\omega)$ together with a Poisson map $J\colon(M,\omega)\to(G,\pi_{0})$ whose image is an open subset of $G$; see [12]. We are interested in the case in which $(M,\omega)$ is exact.
@@
Requiring a symplectic realization $(M,\omega)$ to be exact, $\omega=-\mathrm{d}\theta$, imposes topological restrictions on the realizing manifold $M$ (in particular, $M$ cannot be compact and $\omega]=0$ in de Rham cohomology). This requirement, however, does not constrain the underlying Lie–Poisson or Poisson–Lie structure itself. For a large class of Lie–Poisson systems, including linear and solvable cases, canonical exact realizations are provided by cotangent bundles ${\mathsf{T}}^{\ast}Q$, with the canonical symplectic structure $\omega_{Q}=-\mathrm{d}\theta$ defined by the canonical one-form $\theta_{Q}$ (which is an exact symplectic structure by definition). In that case, the Liouville vector field is the generator of homotheties on the fibers of ${\mathsf{T}}^{\ast}Q$. The example in Section 3.4 illustrates this situation explicitly. More generally, since our construction only requires working on open dense subsets where the homogeneous integrals are independent, the assumption of exactness should be regarded as a natural geometric convenience rather than a substantive restriction of the Poisson–Lie framework.
@@
  • a $1$-homogeneous Hamiltonian $H\in\mathscr{C}^{\infty}(M)$ with respect to the Liouville vector field $\Delta$ of $(M,\theta)$, namely,
    $\Delta H=H\,,\quad\iota_{\Delta}\omega=-\theta\,$,
@@
In many examples of interest, the Poisson pencil $(\pi_{0},\pi_{1})$ admits a symplectic realization $(M,\omega)$ on which the pullbacks of $\pi_{0},\pi_{1}$ give rise to a Poisson–Nijenhuis structure $(\omega,N)$ on $M$ [36, 17, 4, 5]. The eigenvalues of $N$ then provide a natural candidate for homogeneous integrals in involution, and the homogeneity of $H$ can often be arranged by a suitable choice of exact symplectic structure (or by considering an appropriate homogeneous extension).
@@ 3.3 Contact realizations from homogeneous lifts
We first recall a standard fact about exact symplectic manifolds (see, e.g., [39, 13, 43]).
@@
Hence $\Theta\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\theta\wedge(\mathrm{d}\theta)^{n}$ is non-vanishing at every point where $\Delta\neq 0$.
@@
Let $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ and $(M,\theta,H)$ be as in Assumption 3.4. Let $\Delta$ be the Liouville vector field of $(M,\theta)$ and let $C\subset M$ be a hypersurface transverse to $\Delta$, contained in the open subset where $H\neq 0$. Define
@@
By assumption, $H$ is $1$-homogeneous with respect to $\Delta$, i.e., $\mathcal{L}_{\Delta}H=H$. This implies that
@@
where $\varphi_{s}$ denotes the flow of $\Delta$. Hence,
@@
In particular, on the open subset where $H\neq 0$, the level sets of $H$ are transversal to the Liouville flow, and every Liouville orbit intersects each nonzero level set exactly once. By construction $C$ is a hypersurface transverse to $\Delta$ and contained in $\{H\neq 0\}$, so every Liouville orbit in the region of interest intersects $C$ exactly once.
@@
Standard results on the contactization of homogeneous Hamiltonian systems ensure that:
@@
Moreover, since the $F_{j}$ are in Poisson-involution on $M$ and the Jacobi bracket on $C$ is obtained by reduction along the Liouville flow (see [55]), the restrictions $I_{j}$ are in Jacobi-involution:
@@
(ii) Functional independence. Let $U\subset M$ be the dense open subset where $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are independent. Since $C$ is transverse to the Liouville flow, the intersection $U\cap C$ is a dense open subset of $C$. At any $x\in U\cap C$ the differentials $\mathrm{d}F_{0},\dots,\mathrm{d}F_{n}$ are linearly independent in ${\mathsf{T}}^{\ast}_{x}M$. Restricting to ${\mathsf{T}}_{x}C\subset{\mathsf{T}}_{x}M$ we obtain the differentials $\mathrm{d}I_{0},\dots,\mathrm{d}I_{n}$, which remain linearly independent on ${\mathsf{T}}_{x}C$: otherwise there would be a non-trivial linear relation among the $\mathrm{d}F_{j}$ on ${\mathsf{T}}_{x}M$. Thus,
@@
The construction above crucially uses Poisson compatibility and homogeneity on the exact symplectic realization $(M,\theta)$, rather than compatibility of Jacobi structures directly on the contact manifold $(C,\alpha)$. As shown in our previous work on the Jacobi/bi-Hamiltonian no-go phenomenon [10, 43] (building on [53, 47, 50, 9, 52, 55]), compatible Jacobi pairs on a contact manifold do not provide enough independent eigenvalues of the associated recursion operator to recover a full set of contact integrals. By passing through the homogeneous Poisson picture on $(M,\theta)$ we circumvent this obstruction: the bi-Hamiltonian structure lives in the Poisson category, while the contact dynamics is obtained by projection along the Liouville flow. This is summarized in the following diagram:
@@
that, we pass from a bi-Hamiltonian Poisson–Lie system on $(G,\pi_{0},\pi_{1})$ to a contact integrable system $(C,\alpha,h)$ via an exact, homogeneous symplectic realization $(M,\theta,H)$ and a Liouville-transverse hypersurface $C$.
@@
This is an immediate consequence of Proposition 3.6, together with the fact that Liouville-transverse hypersurfaces can be chosen inside any given non-zero level set of $H$ in the region where the homogeneous integrals $(F_{0},\dots,F_{n})$ are independent. ∎
3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two non-trivial contact integrals